NEW DISPERSED FLUORESCENCE SPECTRUM OF THE CBr_{2} Ã–X̃ TRANSITION

HUI-JU HSU, WEI-ZHONG CHANG, and \underline{BOR-CHEN CHANG}, Department of Chemistry, National Central University, 300 Jung-Da Road, Chung-Li 32001, Taiwan.

New dispersed fluorescence spectra following the excitation of the CBr_{2} Ã^{1}B_{1} ←X̃^{1}A_{1} 2_{0}^{13} and 2_{0}^{14} bands at visible wavelengths in a discharge free jet expansion were acquired using an intensified charge-coupled device (ICCD) detector. The new dispersed fluorescence spectra show an improvement of the signal-to-noise ratios by a factor of roughly 10 with respect to our previous report.^{a} These new data extend the scan range from the maximum redshift of 2500 cm^{−1} in our previous report^{a} to 6000 cm^{−1} and reveal more details of the X̃^{1}A_{1} state vibrational structure. Vibrational parameters including fundamental frequencies, anharmonicities, and coupling constants were determined for the CBr_{2} X̃^{1}A_{1} state. Moreover, additional vibrational structures, other than that of the singlet ground electronic state, starting at approximately 2725 cm^{−1} were clearly observed in the new dispersed fluorescence spectra. Since these additional vibrational structures possibly belong to the background triplet state (ã^{3}B_{1}), the singlet-triplet energy gap could be determined. Our progress on the experiments and analyses will be presented.

aC.-L. Lee, M.-L. Liu, and B.-C. Chang, Phys. Chem. Chem. Phys. \underline{5}, 3859-3863 (2003).